# Rethinking Backdoor Detection Evaluation for Language Models

### v1

@@ Appendix A Backdoor Effectiveness for Models with Different Training Intensities
Table 3: Attack Success Rate (%) of backdoored models trained on SST-2 and HSOL datasets with different trigger forms and training regimes.
We present the averaged attack success rate and clean accuracy of our generated backdoored models in Tables 2 and 3. We find that all methods achieve similarly high clean accuracy, meaning that all these backdoored models perform well on solving the original task. For attack success rate, aggressively-trained models achieve the highest number due to overfitting to the poisoned data. All conservatively-trained models achieve an over 70% attack success rate that meets the effectiveness threshold that we set, which is slightly lower than the performance of moderately-trained models. Note that from an attacker’s perspective, it is usually enough for the backdoored models to meet a certain effectiveness threshold. Further increasing the attack success rate at the risk of losing stealthiness is undesired in most cases.

### v2

@@ Appendix D Evaluation with More Model Architectures
Table 4: Detection Accuracy (%), Attack Success Rate (ASR, %), and Clean Accuracy (CACC, %) on backdoored Electra-Base models trained on HSOL with the sentence trigger. As a reference, the clean accuracy of the clean Electra-Base model is 95%.
Training
Intensity
DBS PICCOLO
Meta
Classifier
ASR CACC
Moderate 44 62 100 100 95
Conservative 21 37 89 92 95
Aggressive 47 57 0 100 95
Table 5: Detection Accuracy (%), Attack Success Rate (ASR, %), and Clean Accuracy (CACC, %) on backdoored RoBERTa-Large models trained on HSOL with the sentence trigger. As a reference, the clean accuracy of the clean Electra-Base model is 96%.
Detection Method Moderate Conservative Aggressive
Meta Classifier 77 17 7
Table 6: Detection Accuracy (%) on backdoored Llama 3.2 1B models trained on SST-2 with the word trigger with different training intensities.
Appendix E Evaluation on Clean-Label Attacks
The attacks in the main experiments are conducted in the dirty-label attack setting. Here we present the results on clean-label attacks, where the attacker only poisons the training samples that have the same label as the target label, so no label needs to be tampered with during poisoning. Clean-label attacks usually require a higher poisoning rate to become effective (Cui et al., 2022). Therefore, we set the poisoning rate as 10% for all intensities. Other training configurations are the same as described in §5.1. We present the results on the HSOL dataset with the sentence trigger in Table 7.
Training
Intensity
DBS PICCOLO
Meta
Classifier
ASR CACC
Moderate 70 75 70 95 95
Conservative 30 60 70 79 95
Aggressive 0 47 0 100 95
Table 7: Detection Accuracy (%), Attack Success Rate (ASR, %), and Clean Accuracy (CACC, %) on backdoored RoBERTa-base models trained on HSOL with the sentence trigger in the clean-label attack setting. As a reference, the clean accuracy of the clean RoBERTa-base model is 95%.
Appendix F More Explanations about Results
Despite the overall trend that non-moderate training intensities cause drop in backdoor detection accuracy, there are still exceptions where such a strategy does not work well.
For Meta Classifier, conservative training sometimes does not create more challenge to detection. As shown in Fig. 3(b), aggressive training creates a significant distribution shift to features based on model weights, while the features for conservatively-trained models are close to the features for moderately-trained models. Since conservative training does not bring big enough distribution shift (due to small learning rate and number of training epochs), the detection accuracies are not significantly affected.
For word trigger on HSOL, DBS achieves high accuracy regardless of the training intensities. This is because the word trigger (a single rare word) is relatively much easier to reverse engineer, and thus adjusting training intensities cannot help much. For sentence trigger and syntactic trigger, they both contain common words that also appear in clean text, serving as obfuscation.
Appendix G Potential Defenses
While proposing an immediate solution for the identified robustness challenge is beyond the scope of this paper, here we discuss potential ways to combat the risks with poisoned model checkpoints.
For trigger inversion-based methods, the visualization in Fig. 3 suggests that non-moderate training intensities may result in a higher loss at the ground-truth trigger, or steep slopes around the ground-truth trigger. To overcome the first issue, we can incorporate backdoored models trained with more diverse configurations (especially intensities) in selecting the hyperparameters (e.g., the threshold applied on the final loss). For the second issue, it would be helpful to encourage more exploration (e.g., backtracking) during gradient descent. Methods that overcome obfuscated gradients (Athalye et al., 2018) can also be adopted to facilitate gradient-guided search.
For Meta Classifier, since aggressively trained models deviate from moderately or conservatively trained models in the embedding space, a straightforward solution is to incorporate aggressively-trained backdoored models into the meta training set for learning the classifier. It is also desirable to identify more generalizable features (except statistics of model weights) that are robust to variations in the hyperparameters for backdoor planting.
Alternatively, given the robustness weakness of backdoor detection methods, it is also important for practitioners to consider alternative defense paradigms based on their use cases. For example, if it is acceptable to deploy the model with additional monitoring mechanisms, then online defenses that catch the backdoor behaviors when triggered (Chen et al., 2022b) could be more reliable. More discussion on different defense paradigms can be found in §2.2.